# thanks: These two authors contributed equallythanks: These two authors contributed equally

### v2

@@ Appendix B Secure multi-party computation
We prove the security of our multi-party computation protocol detailed in Section V via a simulation based argument, using the preliminaries of Appendix A. The measurement sequence described in Stage I along with the classical post-processing of Stage II together accomplish the computation of any Boolean function up to quadratic order in the parties’ inputs. We verify this by tracing through the transformation of the graph state |$\ket{G}$ (see Figure 5) in the stabilizer formalism, demonstrating the information that is transmitted along the graph in the form of phase flips conditioned on both the parties’ inputs and their measurements. We establish security by demonstrating that the classically communicated information in the protocol reveals no new information to any of the other parties about their input, other than what they would learn from the final output of the computation. Subsequently, we detail how classical error correction can be utilized to suppress a bit error on the output of the computation. Finally, we discuss how self-testing can be employed to further secure against a possibly corrupt Source.
B.1 Security
@@
Figure 8: Diagram of the security setting for $\mathcal{P}$, as described in Section V. An ideal device (left) publicly broadcasts the function value $f(\tilde{\mathbf{x}},\tilde{\mathbf{y}})$ for private inputs $(\tilde{\mathbf{x}},\tilde{\mathbf{y}})$. A high-level depiction of $\mathcal{P}$ (right), with $\mathbf{m}_{\mathbf{X}}$ denoting the collective local measurement outcomes for party $\mathbf{X}\in\{\mathbf{P},\mathbf{S},\mathbf{R}\}$ and $\mathbf{c}_{\text{tot}}$ denoting all the public communication in the first part of Stage II. The boxes (red, blue, green) indicate local processing by a given party ($\mathbf{P},\mathbf{S},\mathbf{R}$). Solid (dashed) double lines indicate public (private) classical information. Solid single lines indicate distributed quantum information.
@@
We return to the security of our protocol $\mathcal{P}$, introduced in Section V, and present a formal proof. We adopt a simulation based notion of security, in which we present an idealized evaluation of $f$ that receives inputs privately and returns only the correct function evaluation. Figure 8 depicts this security setting in both the ideal and real worlds. We show that all the classical and quantum information held by an adversary in any attack on the real protocol $\mathcal{P}$ can equivalently be simulated from the ideal evaluation of $f$. Hence all the information an adversary extracts about an honest party’s input through $\mathcal{P}$ is already accessible through an ideal evaluation of $f$. In this paper we address two forms of malicious attacks, depicted in Figure 6, where we assume without loss of generality that $\mathbf{P}$ is the party playing honestly.
@@
    (Secure if only $\mathbf{S}$ cheats.) Suppose $\mathbf{P}$ and $\mathbf{R}$ follow $\mathcal{P}$ honestly, but $\mathcal{S}$ potentially deviates. Let $\rho_{\text{cq}}^{\mathbf{S}}(\tilde{\mathbf{x}},\lambda)$ denote the total cq state held by $\mathbf{S}$ at the end of $\mathcal{P}$, given input $\tilde{\mathbf{x}}$ of $\mathbf{P}$ and side information $\lambda$. Then, there exists a simulator $\mathbb{S}$ for $\mathbf{S}$ interacting with the ideal functionality $f_{\text{ideal}}$ that exactly reproduces $\rho_{\text{cq}}^{\mathbf{S}}(\tilde{\mathbf{x}},\lambda)$ for all $\tilde{\mathbf{x}}$ and $\lambda$.
@@
    (Secure if only $\mathbf{R}$ cheats.) Suppose $\mathbf{P}$ and $\mathbf{S}$ follow $\mathcal{P}$ honestly, but $\mathcal{R}$ potentially deviates. Let $\rho_{\text{cq}}^{\mathbf{R}}(\tilde{\mathbf{x}},\tilde{\mathbf{y}})$ denote the total cq state held by $\mathbf{R}$ at the end of $\mathcal{P}$, given inputs $(\tilde{\mathbf{x}},\tilde{\mathbf{y}})$ of $\mathbf{P}$ and $\mathbf{S}$, respectively. Then, there exists a simulator $\mathbb{S}$ for $\mathbf{R}$ interacting with the ideal functionality $f_{\text{ideal}}$ that exactly reproduces $\rho_{\text{cq}}^{\mathbf{R}}(\tilde{\mathbf{x}},\tilde{\mathbf{y}})$ for all $(\tilde{\mathbf{x}},\tilde{\mathbf{y}})$.
@@
The simulation realized in Figure 6(a) is now straightforward. $\mathbf{S}$ generates independently and uniformly random bits
$\{\hat{m}_{i,1},\hat{m}_{i,6},\hat{m}_{i,8},\hat{m}_{i,10},c_{i,A}|i\in\mathcal{A}\}\cup\{\hat{m}_{i,2},\hat{m}_{i,4},\hat{m}_{i,5},\hat{m}_{i,11},c_{i,B}|i\in\mathcal{B}\}\cup\{\Gamma_{P}\}.$ (30)
For each such set of values, $\mathbf{S}$ prepares a quantum states with stabilizers described by Equations (18) and (23). They then operate some quantum instrument $\mathcal{I}$ and generate the variables $\{\hat{c}_{i,B}|i\in\mathcal{A}\}$ and $\{\hat{c}_{i,A}|i\in\mathcal{B}\}$ just as they would when cheating in $\mathcal{P}$, using their prepared quantum states and whatever other side information $\lambda$ they have. What remains is a generation of $\Gamma_{\mathbf{R}}$ that has the same correlations with other variables as in $\mathcal{P}$. To accomplish this, $\mathbf{S}$ first computes the value $\zeta$ in Equation (29) using his sampled variables. He then computes $\tilde{\mathbf{e}}$ and uses this as his input to $f$, thereby obtains $f(\tilde{\mathbf{x}},\tilde{\mathbf{e}})$. From Equation (28), the correct value of $\Gamma_{\mathbf{R}}$ is given by the sum $f(\tilde{x},\tilde{e})+\zeta$.
@@
Hence, the simulation realizing Figure 6(b) is also straightforward. $\mathbf{R}$ first generates independent and uniformly random bits $\{\hat{m}_{i,3},\hat{m}_{i,7},\hat{m}_{i,9},\hat{m}_{i,12}|i=1,\cdots,\mathfrak{R}\}\cup\{\Gamma_{\mathbf{P}}\}$ and prepares the state with the generators in Equation (32). Then for any pair of inputs $(\tilde{\mathbf{x}},\tilde{\mathbf{y}})$ from $\mathbf{P}$ and $\mathbf{S}$ and public evaluation of the function $f(\tilde{\mathbf{x}},\tilde{\mathbf{y}})$, $\mathbf{R}$ computes $\Gamma_{\mathbf{S}}$ from Equation (35).
B.2 Including error correction
@@
$\frac{R}{R_{0}}\geq\left(4(M-1)N\left\lceil\frac{\ln\left((M-1)N/\sqrt{2\epsilon_{f}}\right)}{\left(\tfrac{1}{2}-\epsilon_{*}\right)^{2}}\right\rceil\right)^{-1},$ (41)
@@
where $R_{0}$ is the scheme-dependent average rate to add a new photon to a graph state. For the HoD scheme we propose, $R_{0}=\eta_{e}R_{\text{rep}}$.
B.3 Self-testing of graph states

### v3

@@ Appendix D Secure multi-party computation
We prove the security of our multi-party computation protocol detailed in Sec. III via a simulation based argument, using the preliminaries of Appendix A. The measurement sequence described in Stage I along with the classical post-processing of Stage II together accomplish the computation of any Boolean function up to quadratic order in the parties’ inputs. We verify this by tracing through the transformation of the graph state |$\ket{G}$ (see Fig. 3) in the stabilizer formalism, demonstrating the information that is transmitted along the graph in the form of phase flips conditioned on both the parties’ inputs and their measurements. We establish security by demonstrating that the classically communicated information in the protocol reveals no new information to any of the other parties about their input, other than what they would learn from the final output of the computation. Subsequently, we detail how classical error correction can be utilized to suppress a bit error on the output of the computation. Finally, we discuss how self-testing can be employed to further secure against a possibly corrupt Source.
D.1 Security
@@
Figure 10: Diagram of the security setting for $\mathcal{P}$, as described in Sec. III. An ideal device (left) publicly broadcasts the function value $f(\tilde{\mathbf{x}},\tilde{\mathbf{y}})$ for private inputs $(\tilde{\mathbf{x}},\tilde{\mathbf{y}})$. A high-level depiction of $\mathcal{P}$ (right), with $\mathbf{m}_{\mathbf{X}}$ denoting the collective local measurement outcomes for party $\mathbf{X}\in\{\mathbf{P},\mathbf{S},\mathbf{R}\}$ and $\mathbf{c}_{\text{tot}}$ denoting all the public communication in the first part of Stage II. The boxes (red, blue, green) indicate local processing by a given party ($\mathbf{P},\mathbf{S},\mathbf{R}$). Solid (dashed) double lines indicate public (private) classical information. Solid single lines indicate distributed quantum information.
@@
We return to the security of our protocol $\mathcal{P}$, introduced in Sec. III, and present a formal proof. We adopt a simulation based notion of security, in which we present an idealized evaluation of $f$ that receives inputs privately and returns only the correct function evaluation. Fig. 10 depicts this security setting in both the ideal and real worlds. We show that all the classical and quantum information held by an adversary in any attack on the real protocol $\mathcal{P}$ can equivalently be simulated from the ideal evaluation of $f$. Hence all the information an adversary extracts about an honest party’s input through $\mathcal{P}$ is already accessible through an ideal evaluation of $f$. In this paper we address two forms of malicious attacks, depicted in Fig. 4, where we assume without loss of generality that $\mathbf{P}$ is the party playing honestly.
@@
    (Secure if only $\mathbf{S}$ cheats.) Suppose $\mathbf{P}$ and $\mathbf{R}$ follow $\mathcal{P}$ honestly, but $\mathbf{S}$ potentially deviates. Let $\rho_{\text{cq}}^{\mathbf{S}}(\tilde{\mathbf{x}},\lambda)$ denote the total cq state held by $\mathbf{S}$ at the end of $\mathcal{P}$, given input $\tilde{\mathbf{x}}$ of $\mathbf{P}$ and side information $\lambda$. Then, there exists a simulator $\mathbb{S}$ for $\mathbf{S}$ interacting with the ideal functionality $f_{\text{ideal}}$ that exactly reproduces $\rho_{\text{cq}}^{\mathbf{S}}(\tilde{\mathbf{x}},\lambda)$ for all $\tilde{\mathbf{x}}$ and $\lambda$.
@@
    (Secure if only $\mathbf{R}$ cheats.) Suppose $\mathbf{P}$ and $\mathbf{S}$ follow $\mathcal{P}$ honestly, but $\mathbf{R}$ potentially deviates. Let $\rho_{\text{cq}}^{\mathbf{R}}(\tilde{\mathbf{x}},\tilde{\mathbf{y}})$ denote the total cq state held by $\mathbf{R}$ at the end of $\mathcal{P}$, given inputs $(\tilde{\mathbf{x}},\tilde{\mathbf{y}})$ of $\mathbf{P}$ and $\mathbf{S}$, respectively. Then, there exists a simulator $\mathbb{S}$ for $\mathbf{R}$ interacting with the ideal functionality $f_{\text{ideal}}$ that exactly reproduces $\rho_{\text{cq}}^{\mathbf{R}}(\tilde{\mathbf{x}},\tilde{\mathbf{y}})$ for all $(\tilde{\mathbf{x}},\tilde{\mathbf{y}})$.
@@
The simulation realized in Fig. 4(a) is now straightforward. $\mathbf{S}$ $\mathbb{S}$ generates independently and uniformly random bits
$\{\hat{m}_{i,1},\hat{m}_{i,6},\hat{m}_{i,8},\hat{m}_{i,10},c_{i,A}|i\in\mathcal{A}\}\cup\{\hat{m}_{i,2},\hat{m}_{i,4},\hat{m}_{i,5},\hat{m}_{i,11},c_{i,B}|i\in\mathcal{B}\}\cup\{\Gamma_{\mathbf{P}}\}.$ (60)
For each such set of values, $\mathbb{S}$ prepares quantum states with stabilizers described by Eqs. (48) and (53). It then operate a quantum instrument $\mathcal{I}$, set by the specific attack enacted by $\mathbf{S}$, and generates the variables $\{\hat{c}_{i,B}|i\in\mathcal{A}\}$ and $\{\hat{c}_{i,A}|i\in\mathcal{B}\}$ just as $\mathbf{S}$ would when cheating in $\mathcal{P}$, using their prepared quantum states and whatever other side information $\lambda$ they have. What remains for $\mathbb{S}$ is to generate a $\Gamma_{\mathbf{R}}$ that has the same correlations with other variables in $\mathcal{P}$. To accomplish this, $\mathbb{S}$ first computes the value $\zeta$ in Eq (59) using its sampled variables. It then computes the adversary’s input $\tilde{\mathbf{e}}$ from this information and uses this as its input to $f$, thereby obtaining $f(\tilde{\mathbf{x}},\tilde{\mathbf{e}})$, and allowing $\mathbb{S}$ to return $\rho_{\text{cq}}^{\mathbf{S}}(\tilde{\mathbf{x}},\lambda)$ to the adversary. From Eq (58), the correct value of $\Gamma_{\mathbf{R}}$ is given by the sum $f(\tilde{\mathbf{x}},\tilde{\mathbf{e}})+\zeta$.
@@
Hence, the simulation realizing Fig. 4(b) is also straightforward. $\mathbb{S}$ first generates independent and uniformly random bits $\{\hat{m}_{i,3},\hat{m}_{i,7},\hat{m}_{i,9},\hat{m}_{i,12}|i=1,\cdots,\mathfrak{R}\}\cup\{\Gamma_{\mathbf{P}}\}$ and prepares the state with the generators in Eq (62). Then for any pair of inputs $(\tilde{\mathbf{x}},\tilde{\mathbf{y}})$ from $\mathbf{P}$ and $\mathbf{S}$ and public evaluation of the function $f(\tilde{\mathbf{x}},\tilde{\mathbf{y}})$, $\mathbf{R}$ $\mathbb{S}$ computes $\Gamma_{\mathbf{S}}$ from Eq (65), and returns $\rho_{\text{cq}}^{\textbf{R}}(\tilde{\mathbf{x}},\tilde{\mathbf{y}})$ to the adversary.
D.2 Including error correction
@@
$\frac{R}{R_{0}}\geq\left(6(M-1)^{2}N^{2}\left\lceil\frac{\ln\left((M-1)N/\sqrt{2\epsilon_{f}}\right)}{\left(\tfrac{1}{2}-\epsilon_{*}\right)^{2}}\right\rceil\right)^{-1},$ (71)
@@
where $R_{0}$ is the scheme-dependent average rate to add a new photon to a graph state. For the simpler emit-then-add scheme we propose, employing only destructive photon measurements, $R_{0}=\eta_{e}R_{\text{rep}}$.
D.3 Self-testing of graph states